# Learning from Heterogeneity: A Dynamic Learning Framework for Hypergraphs

### v1

@@ IV-D Loss Function
[…]
@@
where $\mathcal{L}_{recon}$ is defined in Eq. 6. The process of reconstructing the master node using slave nodes can uncover implicit relationships between the master node and others. However, as pointed out above, this process is sensitive to noisy nodes. Over-reliance on reconstruction loss compromises the performance of our model. To enhance the generalization ability, we need to use a weight hyperparameter $\alpha$, to balance the effects of loss of supervised downstream task and reconstruction loss.
IV-E Model Analysis and Discussion
We analyze the computational cost and summarize the advantages of LFH as below:
IV-E1 Computational Cost
In the process of heterogeneous hypergraph generation, the computational cost mainly lies in the reconstruction of the master node from its candidate slave node set (Eq. 4). As it is essential to generate hyperedges of different types for each node, and the computational cost of heterogeneous hypergraph generation is $\emph{O}\left(d|\mathcal{T}_{v}|N\right)$, where $N$ is the number of nodes, $d$ is the dimensionality of node embedding, and $|\mathcal{T}_{v}|$ is the number of node types. Additionally, for the process of dynamic hypergraph learning, we use the incidence matrix H and nodes embedding X to update hyperedges embedding, with the computational cost of $\emph{O}\left(dMN\right)$. Note the number of hyperedges $M$ is proportional to the product of the $N$ and the number of edge types $|\mathcal{T}_{e}|$, thus the computational cost of hyperedges embedding updating is $\emph{O}\left(d|\mathcal{T}_{e}|N^{2}\right)$. Moreover, the computational cost of multi-head attention node embedding updating is $\emph{O}\left(d^{2}|\mathcal{T}_{e}|N\right)$, where $\mathcal{T}_{e}$ is the number of hyperedge types. Hence, the total computational cost of our proposal is $\emph{O}\left(d|\mathcal{T}_{e}|N^{2}+\left(d|\mathcal{T}_{v}|+d^{2}|\mathcal{T}_{e}|\right)N\right)$. The scale of $d$, $\mathcal{T}_{e}$ and $\mathcal{T}_{v}$ are usually far less than $N$, This computational cost is on par with many existing hypergraph-based models, such as [12] and [34].
IV-E2 Discussion
@@ V Experiment
In this section, we empirically evaluate the effectiveness of our LFH framework and analyze the impacts of its key components to the final performance. We conduct our experiments on three different graph datasets, including DBLP, IMDB, and ACM. Additionally, we compare the performance of LFH with eleven baseline models, which include homogenous graph learning, heterogeneous graph learning, and hypergraph learning models.
@@ V-A Experimental Setup
Dataset # Nodes # Node Types # Edges # Edge Types # Node Features
ACM 8994 3 12961 4 1902
DBLP 18405 3 33973 4 334
IMDB 12772 3 18644 4 1256
@@ V-A1 Dataset
  • IMDB333 https://s3.cn-north-1.amazonaws.com.cn/dgl-data/dataset/openhgnn/imdb4GTN.zip contains movie (M), actors (A) and directors (D). Each movie is labelled according to its genre (Action, Comedy, Drama). Node features are also initialized using bag-of-words.
@@ V-A2 Baseline Models and Configurations
We compare with some state-of-the-art baselines, each of which reports promising results in the classification task. Although these baselines share the same goal of learning the representation of the graph, they are originally designed to fulfil graph learning on different graph data, including homogeneous pairwise graph learning, pairwise heterogeneous graph learning, and hypergraph learning.
@@
The shared parameter setup of the framework can be found in Table III. We randomly split all datasets into train/validate/test with the ratio of 0.2/0.1/0.7, respectively. It is worth mentioning that the hyperparameters of all baseline models are set to the optimal values reported in their original papers. All models are trained with a fixed 100 epochs, using an early stopping strategy when the performance on the validation set is not improved for 30 consecutive epochs. All trainable parameters of the neural network are initialized through Xavier [37] and optimized using Adam [38] with the learning rate 2e-3. The dropout rate is set to be 0.3 [39]. The optimal selection of hyperparameters of our model is further analyzed in the following section.
TABLE III: Shared parameter setup
Name  Setup
Optimizer Adam [38]
Learning rate 2e-3
Dropout rate 0.3
Weight initializer Xavier [37]
Epoch 100
MLP 1 linear layer, 1 output layer
@@
Dataset Data split ratio Hypergraph-based Heterogeous-based Homogeous-based
LFH HGNN HGNN+ HL PC-HGN HAN R-HGNN HGT GCN GAT GraphSAGE GraphSAINT
ACM 10% 0.818 0.557 0.655 0.471 0.768 0.551 0.512 0.517 0.505 0.547 0.548 0.512
30% 0.871 0.801 0.851 0.544 0.845 0.842 0.847 0.814 0.762 0.767 0.763 0.771
50% 0.882 0.804 0.852 0.596 0.849 0.842 0.854 0.828 0.846 0.853 0.851 0.839
70% 0.905 0.840 0.882 0.607 0.872 0.866 0.864 0.857 0.853 0.858 0.858 0.856
IMDB 10% 0.491 0.410 0.432 0.415 0.407 0.372 0.406 0.341 0.322 0.384 0.351 0.361
30% 0.539 0.442 0.463 0.424 0.459 0.398 0.415 0.388 0.384 0.429 0.388 0.409
50% 0.558 0.463 0.446 0.437 0.461 0.427 0.438 0.424 0.438 0.458 0.427 0.422
70% 0.571 0.494 0.492 0.459 0.488 0.464 0.487 0.435 0.445 0.488 0.454 0.460
DBLP 10% 0.861 0.669 0.819 0.503 0.818 0.812 0.672 0.598 0.535 0.574 0.543 0.597
30% 0.874 0.739 0.842 0.575 0.838 0.837 0.783 0.632 0.699 0.693 0.696 0.676
50% 0.888 0.767 0.866 0.634 0.861 0.855 0.853 0.724 0.755 0.775 0.771 0.776
70% 0.917 0.791 0.892 0.656 0.884 0.878 0.863 0.782 0.772 0.780 0.778 0.778
@@
V-B Performance Analysis
V-B1 Node Classification
In the node classification task, we split the original data into training data and test data. The training data are then split into a range of [10%, 30%, 50%, 70%]. The strategy of generating the training dataset is the same as [41], where each randomly selected node at least has one pairwise relation with others. The best F1 results among all datasets are demonstrated in Table IV, in which the highest scores are marked in bold. We compared three different categories of graph learning models: homogeneous pairwise graph learning models, heterogeneous pairwise graph learning models, and hypergraph learning models. Our proposed LFH achieves performance gain over other baseline models in all datasets by 2%-35.8%. Our model outperforms all baseline models in all datasets by an average improvement of 12.5%. Compared to four homogeneous pairwise GNN baselines GCN, GAT, GraphSAGE and GraphSAINT, the average F1 score improvement of LFH for all datasets is 14.4%. Regarding the cherry-picked heterogeneous pairwise graph learning models and hypergraph learning models, our model achieved an average improvement of 9.7% and 13.7%, respectively, compared to the best-recorded results in all datasets. It is worth noting that HGNN+ [30] is the only hypergraph learning baseline that utilizes pairwise information for hyperedge generation, resulting in better performance compared to other hypergraph learning models, except for our model. Furthermore, the training process of our model, which incorporates type-specific hyperedge generation and attentive embedding updates that exploits the heterogeneity of the graph, enables LFH to improve HGNN+ by an average of 5.7% in all datasets. Additionally, compared to the latest reporting heterogeneous learning model PC-HGN [7], the best results of our model gains an average increase by 5.2%. For different split ratios of training data, we observe that LFH still outperforms all baselines. When the split ratios are 70%, 50%, 30% and 10%, respectively, our model outperforms other models by 8.6%, 8.7%, 10.9% and 17.7% on average. The obtained result demonstrates that the robustness and generalization capability of our model in the informative case.
V-B2 Link Prediction
@@
To investigate the effectiveness of our proposed framework, we also apply our model to the link prediction task. In this task, a graph with a certain faction of edges removed is given, and thus the objective is to predict these missing edges. Following the experiment setting adopted in [40], 50% percent of existing edges are randomly set hidden as positive samples while ensuring that the residual graph after the edge being removed remains connected. To generate negative examples, we randomly sample an equal number of node pairs from the graph, which have no edge connecting them. Similar to the related works [10, 42], we use a logistic regression classifier as a predictor for link prediction. To get the representation of each edge, we use Hadamard operator to compute the element-wise product for the linked node pairs. We compare the performance of LFH with four baselines. Node2vec [40] is a classic model for link prediction task. The other three baselines, GAT, PC-HGN, and HGNN+, are the top-performing baselines in homogenous pairwise graph learning models, heterogeneous pairwise graph learning models, and hypergraphs learning models in node classification tasks. Experimental results are summarized in Table V. LFH obtains performance gains of 16.8%, 18.8%, 2.9% and 14.6%, compared with Node2vec, GAT, PC-HGN, and HGNN+ and outperforms four baseline models in three datasets by an average improvement of 13.3%. It shows the promising capability of our model to capture implicit data relations. It is worth noting that HGNN+ that performs well in node classification does not have comparable performance in link prediction. This is because the HGNN+ applied for link prediction task heavily relies on pairwise edges to generate hyperedges, and thus a sharp decrease in the number of existing edges leads to a deterioration in performance.
V-C Impact of Hyperedge Construction
In our proposed LFH, $\lambda$ and $\gamma$ are used as hyperparameters. Specifically, $\lambda$ is the weight hyperparameter of the reconstruction error and $\gamma$ is the norm hyperparameter to trade off $l_{1}\text{-}norm$ regularization norm and $l_{2}\text{-}norm$ regularization of the reconstruction coefficient vector (See Eq. 6). Figure 5 and 6 illustrate the overall performance of LFH influenced by $\lambda$ and $\gamma$. Figure 5 illustrates the influence of $\lambda$ for different datasets on F1 score. In this figure, we fix $\gamma$ as 0.02, 0.2 and 0.2 while $\lambda$ ranges between 0.0001 and 1. When $\lambda$ is small, the reconstruction loss of $\lambda c^{k}\left(v^{m}_{i}\right)$ becomes trivial, and the reconstruction coefficient vector p will not experience large adjustments throughout the training phase, thus affecting the generation of hyperedges and deteriorating the performance of LFH. In contrast, when $\lambda$ increases to around 0.2, reconstruction loss of $\lambda c^{k}\left(v^{m}_{i}\right)$ accounts for the appropriate proportion of the total loss, thus achieving the best performance under this setup. When $\lambda$ continually increases to 20, F1 score drops rapidly, which clearly demonstrates that the dominance of reconstruction loss compromises the generalization ability of LFH.
@@ V-C Impact of Hyperedge Construction
V-D Sensitivity Analysis
@@ V-D1 Impact of the node embedding size
We test the impact of different node embedding sizes for all datasets with data split ratio of 10%. As shown in Table VI, the performance of the model increases at first and then starts to drop, reporting the best result with the embedding size 256. Intuitively, larger dimension sizes of node embedding bring in extra redundancies in the training process, thus causing unexpected performance drops.
@@ V-D2 Impact of the number of muti-head
We analyze the effect of the number of multi-head $K$. As observed in Table VII, when the number of attention heads increases, there is typically an improvement in the performance of our model. However, it is observed that the performance of LFH improves only marginally afterwards while causing large computational costs. When $K=4$, it is an optimal point to trade off the performance and computational cost.
@@
Figure 7: Ablation study on pairwise fusion for different models.
@@
V-E Ablation Study
We study the impact of pairwise fusion in this part. We consider different candidates in the pairwise fusion process, aiming to find the best match that is capable of deriving high-quality features from the heterogeneous pairwise graph as the initial node embedding. We analyze the importance of pairwise fusion and implement different models introduced in Section V-A. Note that we applied the same hyperparameters of all these models according to the best performance reported in the respective papers and source codes. As shown in Figure 7, using PC-HGN in pairwise fusion enlarges the performance gap over other models in all datasets, with a performance gain of more than 5.4%-14.4% compared to using GAT, a performance gain of more than 6.6%-18.0% compared to using HGT, and an overall 11.5% performance gain on average compared to using other models in pairwise fusion. This result demonstrates the initial node embedding generated by PC-HGN helps our framework achieve the best results in all three datasets.
V-F Impact of Hyperparameter $\alpha$
As defined in Eq. 6, the proposed united loss is a linear combination of the reconstruction loss in dynamic heterogeneous hyperedges generation and the supervised loss for the downstream task. We further study the impact of $\alpha$ on the performance of LFH. In the experiment, the sensitivity of $\alpha$ is explored to control the level of impact caused by the reconstruction of each master when training the model. When the value of $\alpha$ is close to 1, it means that the model is optimized towards the reconstruction correctness. Fig. 8 reveals the change of F1 score along with $\alpha$. With $\alpha$ ranging in $\left[0.001,0.5\right]$, the performance first reaches to the top and then drops significantly, indicating that over-reliance on reconstruction loss aggravates the generalization ability of our model. We observe that F1 score reaches the highest when $\alpha=0.1$ in all three datasets. The result demonstrates that uniting reconstruction loss in the training process to some extent can help improve the model performance.

### v2

@@ IV-D Loss Function
$\mathcal{L}_{label}=\textit{CE}\left(\textit{MLP}\left(\emph{Z}^{\mathcal{V}}\right),\emph{Y}^{\mathcal{V}}\right)$ (16)
@@
where $\mathcal{L}_{recon}$ is defined in Eq. 6. The process of reconstructing the master node using slave nodes can uncover implicit relationships between the master node and others. However, as pointed out above, this process is sensitive to noisy nodes. Over-reliance on reconstruction loss compromises the performance of our model. To enhance the generalization ability, we need to use a weight hyperparameter $\alpha$, to balance the effects of loss of supervised downstream task and reconstruction loss. The full process of the model, as described in Algorithm 1, and the detailed analysis of computational cost are provided in Appendix.
IV-E Discussion
@@ V Experiment
In this section, We conduct our experiments on three different graph datasets, including DBLP, IMDB, and ACM. We aim to answer the following research questions:
  • RQ1: How does our proposed LFH perform compared with the three kinds of baselines (homogenous graph learning, heterogeneous graph learning, and hypergraph learning models) for node classification task (RQ1.1) and link prediction task (RQ1.2).
  • RQ2: What is the impact of hyperparameters $\lambda$ and $\gamma$ in hyperedge construction on the LFH performance?
  • RQ3: How much do the node embedding size and the number of multi-head $K$ affect the final results?
  • RQ4: How do different components, such as pairwise fusion, dynamic hypergraph construction and multi-head attention mechanism, affect the performance of LFH.
  • RQ5: What is the impact of hyperparameter $\alpha$ on the unified loss?
@@ V-A Experimental Setup
Dataset ACM DBLP IMDB
# Nodes 8994 18405 12772
# Node Types 3 3 3
# Edges 12961 33973 18644
# Edge Types 4 4 4
# Node Features 1902 334 1256
@@ V-A1 Dataset
  • IMDB333 https://s3.cn-north-1.amazonaws.com.cn/dgl-data/dataset/openhgnn/imdb4GTN.zip contains Movie (M), Actor (A) and Director (D). Each movie is labelled according to its genre (Action, Comedy, Drama). Node features are also initialized using bag-of-words.
TABLE III: Shared parameter setup.
Name  Setup
Optimizer Adam [44]
Learning rate 2e-3
Dropout rate 0.3
Weight initializer Xavier [45]
Epoch 100
MLP 1 linear layer, 1 output layer
@@ V-A2 Baseline Models and Configurations
We compare with some state-of-the-art baselines, each of which reports promising results in the classification task. Although these baselines share the same goal of learning the representation of the graph, they were originally designed to fulfil graph learning on different graph data, including homogeneous pairwise graph learning, pairwise heterogeneous graph learning, and hypergraph learning.
@@
  • DHGNN [35] utilizes the k-NN method and k-means clustering to dynamically construct the hypergraph, from which the hypergraph convolution is conducted for node representation learning. The number of clusters and nearest neighbors is set to 64.
The shared parameter setup of the framework can be found in Table III. We randomly split all datasets into train/validate/test with the ratio of 0.2/0.1/0.7, respectively. It is worth mentioning that the hyperparameters of all baseline models are set to the optimal values reported in their original papers. All models are trained with a fixed 100 epochs, using an early stopping strategy when the performance on the validation set is not improved for 30 consecutive epochs. All trainable parameters of the neural network are initialized through Xavier [45] and optimized using Adam [44] with the learning rate 2e-3. The dropout rate is set to be 0.3 [46]. The optimal selection of hyperparameters of our model is further analysed in the following section.
@@
Dataset DSR Hypergraph-based Heterogeneous-based Homogeous-based
LFH HGNN HGNN+ HL DHGNN PC-HCN HAN R-HGNN HGT GCN GAT GraphSAGE GraphSAINT
ACM 10% 0.818 0.557 0.655 0.471 0.486 0.768 0.551 0.512 0.517 0.505 0.547 0.548 0.512
30% 0.871 0.801 0.851 0.544 0.513 0.845 0.842 0.847 0.814 0.762 0.767 0.763 0.771
50% 0.882 0.804 0.852 0.596 0.781 0.849 0.842 0.854 0.828 0.846 0.853 0.851 0.839
70% 0.905 0.840 0.882 0.607 0.881 0.872 0.866 0.864 0.857 0.853 0.858 0.858 0.856
IMDB 10% 0.491 0.410 0.432 0.415 0.417 0.407 0.372 0.406 0.341 0.322 0.384 0.351 0.361
30% 0.539 0.442 0.463 0.424 0.440 0.459 0.398 0.415 0.388 0.384 0.429 0.388 0.409
50% 0.558 0.463 0.446 0.437 0.471 0.461 0.427 0.438 0.424 0.438 0.458 0.427 0.422
70% 0.571 0.494 0.492 0.459 0.528 0.488 0.464 0.487 0.435 0.445 0.488 0.454 0.460
DBLP 10% 0.861 0.669 0.819 0.503 0.481 0.818 0.812 0.672 0.598 0.535 0.574 0.543 0.597
30% 0.874 0.739 0.842 0.575 0.512 0.838 0.837 0.783 0.632 0.699 0.693 0.696 0.676
50% 0.888 0.767 0.866 0.634 0.751 0.861 0.855 0.853 0.724 0.755 0.775 0.771 0.776
70% 0.917 0.791 0.892 0.656 0.784 0.884 0.878 0.863 0.782 0.772 0.780 0.778 0.778
V-B Performance Analysis (RQ1)
V-B1 Node Classification (RQ1.1)
In the node classification task, we initially perform a random sampling of nodes at varying percentages (10%, 30%, 50%, 70%) of the nodes from the three datasets (ACM, IMDB and DBLP). Then, any edges connected to nodes that are not sampled are eliminated. Subsequently, any sampled nodes that lack neighboring nodes are also removed. It is observed that a lower data sampling ratio correlates with an increased introduction of noise into the training dataset, due to the loss of original information. The best F1 results among all datasets are demonstrated in Table IV, in which the highest scores are marked in bold. We compared three different categories of graph learning models: homogeneous pairwise graph learning models, heterogeneous pairwise graph learning models, and hypergraph learning models. Our proposed LFH achieves performance gain over other baseline models in all datasets by 2%-38%. Our model outperforms all baseline models in all datasets by an average improvement of 12.9%. Compared to four homogeneous pairwise GNN baselines GCN, GAT, GraphSAGE and GraphSAINT, the average F1 score improvement of LFH for all datasets is 14.4%. Regarding the cherry-picked heterogeneous pairwise graph learning models and hypergraph learning models, our model achieved an average improvement of 9.7% and 14.6%, respectively, compared to the best-recorded results in all datasets. It is worth noting that HGNN+ [20] is the only hypergraph learning baseline that utilizes pairwise information for hyperedge generation, resulting in better performance compared to other hypergraph learning models, except for our model. Furthermore, the training process of our model, which incorporates type-specific hyperedge generation and attentive embedding updates that exploits the heterogeneity of the graph, enables LFH to improve HGNN+ by an average of 5.7% in all datasets. Additionally, compared to the latest reporting heterogeneous learning model PC-HGN [8], the best results of our model gains an average increase by 5.2%. In general, we observe that our model outperforms all baselines for different data sample ratios (DSR), which is shown in Table IV. When the data sample ratios are set to be 70%, 50%, 30% and 10%, our model outperforms other models by 10.1%, 10.2%, 10.9% and 16.7% on average, respectively. It is worth noting that our model achieves higher average improvements when the data sampling ratio is set to 10%, and then the margin of average improvements declines as the data sampling ratio increases.
@@ V-B Performance Analysis (RQ1)
V-B2 Link Prediction (RQ1.2)
@@
To investigate the effectiveness of our proposed framework, we also apply our model to the link prediction task. In this task, a graph with a certain faction of edges removed is given, and thus the objective is to predict these missing edges. Following the experiment setting adopted in [47], 50% percent of existing edges are randomly set hidden as positive samples while ensuring that the residual graph after the edge being removed remains connected. To generate negative examples, we randomly sample an equal number of node pairs from the graph, which have no edge connecting them. Similar to the related works [16, 48], we use a logistic regression classifier as a predictor for link prediction. To get the representation of each edge, we use Hadamard operator to compute the element-wise product for the linked node pairs. We compare the performance of LFH with five baselines. Node2vec [47] is a classic model for link prediction task. DHGNN is a hypergraph learning method that incorporates dynamic hypergraph construction. The other three baselines, GAT, PC-HGN, and HGNN+, are the top-performing baselines in homogenous pairwise graph learning models, heterogeneous pairwise graph learning models, and hypergraphs learning models in node classification tasks. Experimental results are summarized in Table V. LFH obtains performance gains of 16.8%, 18.8%, 2.9%, 14.6% and 13.7%, compared with Node2vec, GAT, PC-HGN, HGNN+, DHGNN and outperforms five baseline models in three datasets by an average improvement of 12.8%. It shows the promising capability of our model to capture implicit data relations. It is worth noting that HGNN+ that performs well in node classification does not have comparable performance in link prediction. This is because the HGNN+ applied for link prediction task heavily relies on pairwise edges to generate hyperedges, and thus a sharp decrease in the number of existing edges leads to a deterioration in performance.
@@
V-C Impact of Hyperedge Construction (RQ2)
In our proposed LFH, $\lambda$ and $\gamma$ are used as hyperparameters. Specifically, $\lambda$ is the weight hyperparameter of the reconstruction error and $\gamma$ is the norm hyperparameter to trade off $l_{1}\text{-}norm$ regularization norm and $l_{2}\text{-}norm$ regularization of the reconstruction coefficient vector (See Eq. 6). Figure 4 and 5 illustrate the overall performance of LFH influenced by $\lambda$ and $\gamma$. Figure 4 illustrates the influence of $\lambda$ for different datasets on F1 score. In this figure, we fix $\gamma$ as 0.02, 0.2 and 0.2 while $\lambda$ ranges between 0.0001 and 1. When $\lambda$ is small, the reconstruction loss of $\lambda c^{k}\left(v^{m}_{i}\right)$ becomes trivial, and the reconstruction coefficient vector p will not experience large adjustments throughout the training phase, thus affecting the generation of hyperedges and deteriorating the performance of LFH. In contrast, when $\lambda$ increases to around 0.2, reconstruction loss of $\lambda c^{k}\left(v^{m}_{i}\right)$ accounts for the appropriate proportion of the total loss, thus achieving the best performance under this setup. When $\lambda$ continually increases to 20, F1 score drops rapidly, which clearly demonstrates that the dominance of reconstruction loss compromises the generalization ability of LFH.
V-D Sensitivity Analysis (RQ3)
@@ V-D1 Impact of the node embedding size
We test the impact of different node embedding sizes for all datasets with data split ratio of 10%. As shown in Table VI, the performance of the model increases at first and then starts to drop, reporting the best result with the embedding size 256. Intuitively, larger embedding sizes could lead to overfitting, thus causing unexpected performance drops.
@@ V-D2 Impact of the number of muti-head
We analyse the effect of the number of multi-head $K$. As observed in Table VII, when the number of attention heads increases, there is typically an improvement in the performance of our model. However, it is observed that the performance of LFH improves only marginally afterwards while causing large computational costs. When $K=4$, it is an optimal point to trade off the performance and computational cost.
@@
Figure 6: Ablation on using different GNN models for initial embedding generation
TABLE VIII: Performance of LFH for ablation study.
Model ACM DBLP IMDB
LFH 0.818 0.861 0.491
-w/o Dynamic Hypergraph Constrcution 0.774 0.673 0.405
-w/o Multi-head Attention 0.653 0.582 0.344
V-E Ablation Study (RQ4)
We study the impact of pairwise fusion in this part. We consider different candidates in the pairwise fusion process, aiming to find the best match that is capable of deriving high-quality features from the heterogeneous pairwise graph as the initial node embedding. We analyse the importance of pairwise fusion and implement different models introduced in Section V-A. Note that we applied the same hyperparameters of all these models according to the best performance reported in the respective papers and source codes. As shown in Figure 6, using PC-HGN in pairwise fusion enlarges the performance gap over other models in all datasets, with a performance gain of more than 5.4%-14.4% compared to using GAT, a performance gain of more than 6.6%-18.0% compared to using HGT, and an overall 11.5% performance gain on average compared to using other models in pairwise fusion. This result demonstrates the initial node embedding generated by PC-HGN helps our framework achieve the best results in all three datasets.
In order to validate the effectiveness of different components in LFH, we design two variants of LFH, and compare the performance of these variants with LFH in the node classification task. The results of comparison are shown in Table VIII. Specifically, we investigate the contributions of dynamic hypergraph construction and multi-head attention to the performance of LFH. LFH w/o dynamic hypergraph construction replaces the learning process with unsupervised $k$-NN method for generating heterogeneous hyperedges, and the number of nearest nodes to form a hyperedge is set to 10 [19]. The results show that LFH has a better performance over LFH w/o dynamic hypergraph construction by margins of 4.4%-18.9%, which indicates that the learning process for generating hypergraph enables LFH to encode topological structure more accurately. LFH w/o multi-head attention excludes the importance of heterogeneous hyperedges with respect to the master node, and assigns the weights to different heterogeneous hyperedges equally during hypergraph learning. The performance of LFH w/o multi-head attention drops significantly in all three datasets. The result demonstrates the significance of the dynamically learned weights of hyperedges and the effectiveness of the multi-head attention mechanism.
@@ V-E Ablation Study (RQ4)
V-F Impact of Hyperparameter $\alpha$ (RQ5)
As defined in Eq. 6, the proposed united loss is a linear combination of the reconstruction loss in dynamic heterogeneous hyperedges generation and the supervised loss for the downstream task. We further study the impact of $\alpha$ on the performance of LFH. In the experiment, the sensitivity of $\alpha$ is explored to control the level of impact caused by the reconstruction of each master when training the model. When the value of $\alpha$ is close to 1, it means that the model is optimised towards the reconstruction correctness. Fig. 7 reveals the change of F1 score along with $\alpha$. With $\alpha$ ranging in $\left[0.001,0.5\right]$, the performance first reaches to the top and then drops significantly, indicating that over-reliance on reconstruction loss aggravates the generalization ability of our model. We observe that F1 score reaches the highest when $\alpha=0.1$ in all three datasets. The result demonstrates that uniting reconstruction loss in the training process, to some extent, can help improve the model performance.
Figure 8: A t-SNE visualization of our proposed model compared to baseline models on the ACM dataset (red indicates paper nodes, green denotes author nodes, and blue represents subject nodes).
V-G Visualization
To investigate the effectiveness of our model compared to other graph learning models in encoding the heterogeneity of the topology, we visualize the learned node embeddings using the t-SNE tool [49] on the ACM dataset. It is worth noting that ACM dataset naturally possess heterogeneous properties, as represented by various types of nodes (such as Author, Paper and Subject in ACM). Visualizing the distributions of nodes with different types directly demonstrates the method’s ability to capture heterogeneity. We compare LFH with other three baselines, GAT, PC-HGN, and HGNN+. These models represent the top-performing baselines for node classification tasks in homogeneous pairwise graph learning, heterogeneous pairwise graph learning, and hypergraph learning. As shown in Fig. 8, our proposed model LFH yields discernible clusters compared to other baseline models on the ACM dataset, which verifies the effectiveness of the proposed method in capturing the heterogeneity of the topology.